OpenStudy (anonymous):

An ice cream cone has a diameter of 2 inches and a height of 4 inches. Find its volume. Round to the nearest tenth.

OpenStudy (anonymous):

do you know the formula for the volume of a cone?

OpenStudy (anonymous):

4.18879

OpenStudy (anonymous):

V=(1/3)r^2.h =(1/3) * 1*4= 4/3

OpenStudy (anonymous):

=1.3

OpenStudy (anonymous):

πr2∗4 v=------ 3

OpenStudy (anonymous):

r^2 sorry

OpenStudy (anonymous):

V=(1/3)pi *r^2.h =(1/3) * 1*4*3.14= 4.1

OpenStudy (anonymous):

I looked up the formula for the volume i think your right raheen

OpenStudy (anonymous):

o got the same thing but he didnt round properly it would be 4.2 because its really 4.18879

OpenStudy (anonymous):

yes it's ok as V=(1/3) pi*r^2.h

OpenStudy (anonymous):

and formulas are written differently it depends on how u like it

OpenStudy (anonymous):

oh, i didnt know. So rounded to the nearest tenth would be 4.2